A company purchased 500 units for $30 each on January 31. It purchased 550 units for $33 each on February 28. It sold a total of 650 units for $45 each from March 1 through December 31. What is the cost of ending inventory on December 31 if the company uses the weighted-average inventory costing method? (Assume that the company uses a perpetual inventory system.)

A) $13,200

B) $12,629

C) $12,000

D) $11,378
Answer #1

Weighted average cost per unit = ( 500 x 30) + (550 x 33)/( 500+550)

= (15,000+18,150)/1,050

= 33,150/1,050

= $31.5714286

Ending inventory = Number of units available for sale - Number of units sold

= 1,050-650

= 400

Cost of ending inventory = Ending inventory x Weighted average cost per unit

= 400 x 31.5714286

= $12,629

Correct option is B.

Kindly comment if you need further assistance.

Thanks‼!